OpenStudy (anonymous):

Could someone check my work? Solve (x + 3)^2 = 36 for x

OpenStudy (anonymous):

√(x+3)^2 = √36 x+3 = 6 -3 -3 x=6

jimthompson5910 (jim_thompson5910):

when you take the square root of both sides, you will have a plus/minus

jimthompson5910 (jim_thompson5910):

for example x^2 = 1 x = +-sqrt(1) x = 1 or x = -1

OpenStudy (anonymous):

x=3

jimthompson5910 (jim_thompson5910):

so your step 2 should be x+3 = 6 or x+3 = -6

OpenStudy (anonymous):

so 3 and -9

OpenStudy (anonymous):

Ok, got cha